OpenStudy (anonymous):

AB and AD are tangents of the circle with the center at C. The measure of BDC = 45o, and the circle has a diameter of 4. Which is the length of AB?

ganeshie8 (ganeshie8):

draw it out

OpenStudy (anonymous):

so how do you do that?

ganeshie8 (ganeshie8):

use this : base angles of isosceles triangle congruent

OpenStudy (anonymous):

what is that?

ganeshie8 (ganeshie8):

lookup isosceles triangle and get some familiarity with that first, then oly you can solve this

OpenStudy (phi):

notice line segment CD and line segment CB are both a radius= 1/2 the diameter = 1/2 * 4 = 2

OpenStudy (phi):

angle CDB is 45º. what is angle CBD ?

OpenStudy (anonymous):

45 as well

OpenStudy (phi):

if you have a triangle with 45, 45, what is the 3rd angle BCD ?

OpenStudy (phi):

knowing all 3 angles add up to 180

OpenStudy (phi):

next, what is the angle CDA formed by the tangent line ?